OpenStudy (anonymous):

Solve for x: 8x + 17 = 5(x + 1)

OpenStudy (lalaly):

why does it all show on one line? :S

OpenStudy (anonymous):

because the line breaks are broken.

OpenStudy (anonymous):

use latex

OpenStudy (lalaly):

oh okayy

OpenStudy (lalaly):

\[8x+17 =5x +5\]\[3x=-12\] \[x=-4\]

OpenStudy (anonymous):

Solve for x: -2(x - 3) = 10

OpenStudy (anonymous):

\(8x+17 = 5(x+1)\)\[\implies 8x + 17 = 5x + 5\]\[\implies 3x = -12\]\[\implies x = -\frac{12}{3} = -4\]

OpenStudy (anonymous):

Ok, that is the same problem. Just do it the same way SaFar.

OpenStudy (anonymous):

Different numbers, but the same deal. Distribute the -2, isolate the variable term, divide by the coefficient.

OpenStudy (anonymous):

-2-6=10 +6 +6 -2 = 16 --- --- 2 2 x= 8

OpenStudy (anonymous):

Why is it on one line?

OpenStudy (anonymous):

It should be -2x + 6

OpenStudy (anonymous):

-2(x-6) = -2(x) + (-2)(-6) = -2x + 6

OpenStudy (anonymous):

x = 8?

OpenStudy (anonymous):

\[-2(x-3)\]\[=-2x + -2(-3)\]\[=-2x + 6\]

OpenStudy (anonymous):

So you mixed up your signs when you distributed. Try it again.

OpenStudy (anonymous):

What i did was -2 + 6 = 10 ----- Equation after being distributed -6 -6 -2 = 4 Divide by 2 x = -2?

OpenStudy (anonymous):

OMG its not looking how i typed it

OpenStudy (anonymous):

x = -2?

OpenStudy (anonymous):

\[-2x + 6 = 10\]\[\implies -2x = 4\]\[\implies x = -\frac{4}{2} = -2\]

OpenStudy (anonymous):

Nicely done

OpenStudy (anonymous):

So when i have -2(x-3) Distribute -2x - 6 = 10 or -2x + 6 = 10?

OpenStudy (anonymous):

Addition or Subtraction sign?